OpenStudy (twizttiez):

Find the slope y-2=-3(x+1) Help very much appreciated

OpenStudy (vjt):

what is your first step?

OpenStudy (twizttiez):

Thats the problem idk :/

OpenStudy (vjt):

did you try distributive property?

OpenStudy (twizttiez):

I missed a day of school and I'm behind

OpenStudy (vjt):

ok ill try to help, but it is against the code of conduct to just give you the answer.

OpenStudy (twizttiez):

I know

OpenStudy (anonymous):

@VJT no offence but for real

OpenStudy (vjt):

what?

OpenStudy (twizttiez):

So what do i do first?

OpenStudy (vjt):

y-2=-3(x+1) you are multiplying -3*x first=-3x then multiply -3*1=-3 y-2= -3x-3

OpenStudy (vjt):

theirs your first step

OpenStudy (anonymous):

nvm

OpenStudy (vjt):

next step is to get y all by itself. y-2= -3x-3 +2 +2 so your adding 2 on both sides of the equation to get... y=-3x-1

OpenStudy (vjt):

that is your answer

OpenStudy (vjt):

the slope would be -3x

OpenStudy (twizttiez):

Ok I'm starting to understand thanks!

OpenStudy (twizttiez):

Would you like a medal?

OpenStudy (vjt):

no problem do you have anymore. i can help. and i kinda wouldn't mind a medal

OpenStudy (vjt):

thanks

OpenStudy (anonymous):

good job @VJT

OpenStudy (twizttiez):

Yes i do

OpenStudy (twizttiez):

y+6=1/3(x-4)

OpenStudy (vjt):

ok can you try to start this one out on your own? or should i show you the first step

OpenStudy (anonymous):

i am not smart

OpenStudy (twizttiez):

Wouldn't it be the same thing?

OpenStudy (vjt):

yes, doing the distributive property first.

OpenStudy (vjt):

y+6=1/3(x-4) 1/3*x=1/3x 1/3*4=1.3

OpenStudy (twizttiez):

So 1.3 is my slope correct?

OpenStudy (anonymous):

correct